# Constructing equivalences between fusion categories of quantum groups and of vertex operator algebras via quantum gauge groups

### v4

@@ 1. Introduction
In two or three dimensional quantum field theory, it was noted that the corresponding categories were braided rather than symmetric. This implies that Bose-Fermi alternative does not hold any longer. Models were first described by Streater and Wilde. Doplicher, Haag, Roberts theory of superselection sectors was generalized to braid groups statistics by Fredenhagen, Rheren, Schroer [35], [36]. The representation of the braid group in the tensor $C^{*}$-category is unitary.
The discovery of quantum groups first arose in physics literature in the work by Faddeev, introduced as noncommutative function algebras. In the mid 80s breakthrough papers by Drinfeld and Jimbo discovered them as quantization of universal enveloping algebras, and developed their properties. We refer to the introduction in [11] and more historical information at the end of each chapter in their book. Woronowicz developed a theory of compact quantum groups as noncommutative function algebras in the setting of $C^{*}$-algebras and constructed the first important examples as deformation of ${\rm SU}(d)$ by a real parameter, see [80] and reference to the original papers, via a concrete Tannaka-Krein duality for $C^{*}$-algebras extending the classical case.
After the discovery of the quantum groups in the mid eighties, Doplicher asked the author to try to extend their theory and consequent applications on the construction of a quantum gauge group and a related field algebra. Of course, several studies were dedicated to extend Doplicher-Roberts duality theory to quantum groups, see Ch. IV.5 in [48] for more references on an early stage of this problem, see also the introduction in [13].
Our approach may be regarded in the setting of Connes noncommutative geometry [15], and follows naturally our previous work in the setting of compact quantum groups, see also the introduction in [14], [16]. In [83], [84], [88], [85], [86], [87], [89] we have developed at the level of rigid tensor $C^{*}$-categories with simple unit an analogue of Doplicher-Roberts construction of a compact gauge group and field algebra, or more precisely the analogue of [23] in the setting of compact quantum groups. The analogy cannot be performed in a strict way due to new difficulties in the quantum case, see the mentioned papers and [13]. One important difference was that for rigid braided tensor categories we do not have braided tensor functors to the categories of vector spaces, beyond the symmetric ones, because the category of vector spaces does not admit a braided symmetry which is not a permutation symmetry. Giving up a compatibility property in the symmetric case implies lack of uniqueness of the functor, and thus of the gauge group. This was noticed early as an easy consequence from their construction.
@@
To describe our approach, we recall that Drinfeld original motivation for introducing quasi-Hopf algebras was to interpret them as symmetries to describe braided tensor categories from important models of rational conformal field theories motivated by the work by Moore and Seiberg. Although as noted in [26], they could not describe precisely simple models such as the Ising, Drinfeld evaluated quasi-Hopf algebras a good description of quantum symmetries for the WZW model in an approximate way, i.e. in the limit to infinity of the central charge.
@@
From this perspective, we regard the equivalence described by Drinfeld-Kohno theorem as a Tannakian counterpart at the level of quasi-Hopf algebras of Kazhdan-Lusztig-Finkelberg equivalence $\tilde{\mathcal{O}}_{k}\to{\mathcal{C}}({\mathfrak{g}},q)$. This point of view is the author understanding of the simplified proof given by Neshveyev and Tuset of Drinfeld-Kohno theorem following the proof by Kazhdan and Lusztig of the same theorem over the complex numbers, and for real values of the deformation parameter $q$, in the setting of compact quantum groups. Their approach inspired us to a similar strategy in our case of an abstract duality theory for both the fusion categories of quantum groups at roots of unity and of affine vertex operator algebras at positive integer levels [78].
There have been studies that construct analogues of Doplicher-Roberts field algebra using generalizations of quasi-Hopf algebras (the weak quasi-Hopf algebras) as quantum gauge groups, and they showed complications reflecting non-triviality of the associator, see [93]. It would be desirable to be able to use quantum symmetry groups that may be regarded in the family of Hopf algebras.
To treat the difficulty of finding a replacement of Doplicher-Roberts symmetric functors, in [12] and [13] we have taken into account Drinfeld notion of coboundary symmetry $\overline{R}$, which plays a role in the construction of Drinfeld twist in Drinfeld-Kohno theorem [27]. This is a symmetry of the category obtained modifying the braided symmetry defined by the $R$-matrix by a square root of the $2$-coboundary defined by the ribbon structure, and by construction the coboundary symmetry is an involutive symmetry, similarly to the permutation symmetry. This symmetry finally allows for a generalization of the notion of symmetric functor. Moreover we have followed both the original choice by Mack and Schomerus [71] of a given dimension function corresponding to non-negligible tensor powers of the vector representation of quantum ${\mathfrak{s}l}_{2}$ at roots of unity and also the study of classification of compact quantum groups corresponding to a given representation category of deformation compact quantum groups by Neshveyev and Yamashita [81], [82]. They both considered a specific positive integer valued dimension function associated to the given category of the quantum group, which was the classical dimension function.
This suggested to us to consider in [12] and [13] a specific dimension function on a functor related to classical data, and this was first Wenzl functor [99] for the quantum group fusion category now at roots of unity, and then the minimum energy functor for the WZW model [106] in the setting of vertex operator algebras. It was particularly important for us that in [99] Wenzl describes an explicit quantization curve associated to his functor that leads to an algebra that identifies with the Zhu algebra for that model in the setting of vertex operator algebras. Since in [12] for the type $A$ case we had constructed a natural quantum group structure on a semisimple associative algebra from the quantum group fusion category endowed with Wenzl functor, compatible with the tensor structure, this arose the question of constructing another natural quantum group structure on the Zhu algebra at the time of [12], following the quantization curve on the unitary structure of tensor products with the aim of making it classical. This involved in [13] the construction of a Drinfeld twist in our setting, the above mentioned notion that Drinfeld had introduced to study equivalence of the Drinfeld category with the quantum group category in a formal setting. We have done this for all the Lie types via our version of Drinfeld-Kohno Theorem 30.4 in [13] and applications to affine vertex operator algebras at positive integer level, via Zhu linear equivalence, in [13].
We refer to Sect. 25 in [13] for an extension of certain results by Neshveyev and Yamashita in the setting of quantum group fusion categories of type $A$ based on Kazhdan-Wenzl theory at roots of unity in our case, and on the weak Hopf algebras constructed in [12], in place of the compact quantum groups. This section is important for us as it suggested that in general the associativity morphisms in a category arising from CFT may be of a simple nature from the viewpoint of the associated quantum gauge groups that need to be constructed, and this finally proved to be the case, we shall come back to this with more precision.
On the other hand, in the process of extending Doplicher-Roberts theory in the setting of compact quantum groups, we had started by embedding in a tensorial way part of the morphisms of the category into the Hilbert spaces. This was the tensor subcategory generated by solutions of the conjugate equations [83], [85], [89]. Via Tannaka-Krein-Woronowicz duality this subcategory corresponds to Wang and Van Daele orthogonal or universal compact quantum groups $A_{o}(F)$, or $A_{u}(F)$ [97], their representation category has been studied by Banica [2], [3].
The choice of the compact quantum group arising in this way depends on whether a generating object $\rho$ is selfconjugate as in the case of $A_{o}(F)$ or not as for $A_{u}(F)$. In doing this embedding we had already made a choice of a dimension function, the rank of the associated invertible matrix $F$, and from that choice we had constructed a tensor functor from that subcategory. To consider the full category, the analogue Doplicher-Roberts compact gauge group is described by an ergodic action of the corresponding quantum group on a quantum homogeneous space rather than a quantum group itself, due to lack of commutativity of this $C^{*}$-algebra [85]. The homogeneous space is described by a $C^{*}$-algebra, naturally associated to the category and to the embedding of the subcategory.
This property is likely to hold also in other models arising from conformal field theory starting from the minimum energy functor, unless one can show that certain known obstructions already met in [84] for compact quantum groups, vanish, we shall come back to this in the outlook section 13.
However, for the WZW models arising from affine Lie algebras at positive integer levels in [12] for the type $A$ case, and in [13] for all the Lie types, we obtain genuine quantum groups as quantum gauge groups, described as semisimple weak Hopf algebras with $R$-matrix, $C^{*}$-structure explicitly associated to Drinfeld coboundary matrix, and with unitary representations of the braid group, derived from the quantum group fusion category. We prove that these weak Hopf algebras have representation category equivalent to both the starting fusion category for all the Lie types and also to that of the corresponding affine vertex operator algebra with ribbon braided Huang-Lepowsky tensor structure, for the classical Lie types and $G_{2}$. Constructing the equivalence from the former to the latter in the setting of semisimple weak quasi-Hopf algebras as we do in [13] for these Lie types is our proof of Kazhdan-Lusztig-Finkelberg theorem, see Theorem 2.1.
Among our first references in the context of conformal field theory important for us to start to set up the connection with the work of Wenzl in quantum groups, is the work by Wassermann [98] in the setting of affine Lie algebras, where the key concept is played by the notion of primary field, and its initial term, which we use in a fundamental way to connect quantum group fusion category to representations of the Zhu algebra and their special tensor products with the fundamental representation, see Sect. 35 in [13]. Most importantly, we describe a connection between Wenzl and Wassermann work via Wenzl quantisation curve, and building on their work in [13] we obtain Theorems 3.1, 3.2, Cor. 3.3, and Theorem 3.4 which compares with the tensor product structure by Huang and Lepowsky, to which we refer the reader also for important historical references. The connection from affine Lie algebra setting to Huang-Lepowsky tensor product theory may be found in Sections 34–39 in [13].
Our abstract duality methods need new further tools as we shall see from the theory of quantum groups, mainly from the work of Drinfeld and Wenzl, that are not available in the theory of groups. However, we shall find that in both contexts, quantum groups at roots of unity and affine vertex operator algebras at positive integer levels, both representation categories are described by functors extending the symmetric tensor functors in Doplicher-Roberts abstract duality theory, see Sect. 29 in [13]. We call these structures hermitian coboundary weak quasi-Hopf algebras, as the hermitian structure is given by Drinfeld coboundary matrix, which is a modification of the $R$-matrix with a square root of the ribbon structure.
We notice on the other hand that a stronger property is needed to establish the connection, than the purely categorical characterisation of functors in Sect. 29 in [13]. This property is expressed in terms of quantum groups, and the first example, in a non-semisimple setting, is the hermitian structure of $U_{q}({\mathfrak{g}})$ at a root of unity by Theorem 31.9 in [13]. This property says that the -involution enjoys a precise anticommutation relation (rather than commutation relation as is familiar in Hopf $C^{*}$-algebras) with the coproduct and was first noticed in [99]. We axiomatize this property in Sect. 30 in [13], as a compatibility property of hermitian coboundary weak quasi-Hopf algebras. Further examples are our weak bi-algebras $A_{W}({\mathfrak{g}},q)$ associated to Wenzl functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$ and to the quasi-bialgebras structures on the Zhu algebra $A({V_{{\mathfrak{g}}_{k}}})$ associated to Zhu functor ${\rm Rep}({V_{{\mathfrak{g}}_{k}}})\to{\rm Vec}$ via our Drinfeld twist construction, see Sections 32, 34 in [13]. The anticommutation property of the -involution and coproduct of $A_{W}({\mathfrak{g}},q)$ is not a direct consequence of the corresponding property of $U_{q}({\mathfrak{g}})$, as its coproduct is not compatible with the fusion tensor product. We need some delicate analysis of the fusion tensor product described in [99]. For $A({V_{{\mathfrak{g}}_{k}}})$, we show that the twist makes the coproduct of $A({V_{{\mathfrak{g}}_{k}}})$ symmetric, and this is a result analogous to the symmetry property of the usual coproduct of Drinfeld quasi-Hopf algebra $U({\mathfrak{g}})[[h]]$. For details we refer to the cited sections.
On the other hand, as an algebra $A_{W}({\mathfrak{g}},q)$ is a natural quotient of $U_{q}({\mathfrak{g}})$, and the antipode of $A_{W}({\mathfrak{g}},q)$ is directly derived from that of $U_{q}({\mathfrak{g}})$. In this way, $A_{W}({\mathfrak{g}},q)$ becomes a weak Hopf algebra. It follows that ${\rm Rep}(A_{W}({\mathfrak{g}},q))$ becomes a rigid ribbon braided tensor category naturally equivalent to ${\rm Rep}(A({V_{{\mathfrak{g}}_{k}}}))$ via our twist, and therefore so does ${\rm Rep}({V_{{\mathfrak{g}}_{k}}})$ via Zhu equivalence.
As above motivated, we shall derive more structure beyond rigidity and modularity on ${\rm Rep}({V_{{\mathfrak{g}}_{k}}})$ from the quantum group fusion category, a further example is unitarization of the vertex operator algebra representation category from the corresponding quantum group fusion category, via the explicit equivalence, see Sects. 22, 23, 24, 26 in [13]. To this aim it is important that our equivalence passes through the weak quasi-Hopf algebras, and the fact that these are naturally associated to the categories. Indeed it was known that a mere tensor equivalence between the categories would not suffice to transport the unitary structures see [49]. Sect. 16 in [13] treats this for general tensor categories that admit submultiplicative dimension functions (called weak dimension functions).
Our equivalence, since it is given by a Drinfeld twist analogous to [25], [27], and is also motivated by the requirement of making the unitary structure classical, at the level of the associated weak Hopf or quasi-Hopf algebras gives an understanding of the unitary structures of representation categories of affine vertex operator algebras as classical compact real forms, or untwisted. We refer to the outlook section for some more information on possible extension of our viewpoint to models beyond the affine vertex operator algebras at positive integer levels, see also [13].
Our main result is Theorem 2.1. In the subsequent sections we outline the plan of our proof. A main idea that we follow is to consider the fundamental representation $V$ of ${\mathfrak{g}}$ for each Lie type as in [99] where the author explains its generating property and good fusion structure with respect to tensor products with an arbitrary irreducible representation allowing a well-defined truncation procedure to obtain non-negligible objects, with the aim of studying unitarizability of the fusion category. It should also be noted that a main application of that paper is construction of inclusion of subfactors of finite Jones index [63] for all Lie types, extending his previous work.

### v5

@@ 1.4. Quantum symmetries and primary field reconstruction in low-dimensional AQFT
In low-dimensional quantum field theories (specifically in two and three spacetime dimensions), the categories of superselection sectors naturally admit a braided tensor structure rather than a symmetric one. Consequently, the standard DHR Bose-Fermi alternative does not hold any longer, allowing for the existence of non-integer statistics. Early models exhibiting such behavior were first constructed by Streater and Wilde. The DHR theory of superselection sectors was subsequently generalized to accommodate braid group statistics by Fredenhagen, Rehren, and Schroer [32, 33], who showed that the representations of the braid group within these tensor $C^{*}$-categories are unitary.
Motivated by the Drinfeld-Kohno theorem and the representation theory of quantum groups at roots of unity–where Hermitian representations of $U_{q}(\mathfrak{sl}_{2})$ become non-semisimple and exhibit null vectors–Mack and Schomerus sought to recover the truncated fusion tensor products of conformal field theory (CFT) from a semisimple quotient algebra. To this end, they extended Drinfeld’s framework to define weak quasi-Hopf algebras. They proposed that the true global quantum symmetries of minimal conformal models are weak quasi-Hopf algebras canonically associated with $U_{q}(\mathfrak{sl}_{2})$ at roots of unity. As a first main example, they rigorously showed that the primary field operators of the chiral Ising model satisfy weak quasi-Hopf-covariant operator product expansions and obey local braid relations [69]. Their program culminated in Schomerus’s generalized construction of a weak quasi-Hopf algebra and a corresponding field algebra for any rational conformal field theory, extending the Doplicher-Roberts reconstruction theorem to the braided case [70], [93].
Mack-Schomerus quantum symmetry program saw further but limited development. A primary obstacle was that the emerging quantum gauge groups in the general theory were highly non-unique and possessed a complicated algebraic structure, largely due to the non-triviality of the associator. These ambiguities remained unclarified, leaving an incomplete reconciliation with the standard Doplicher-Roberts theory. Furthermore, the pioneering examples provided by Mack and Schomerus were derived from quantum group representation theory, rather than reconstructed intrinsically from the local observables. This approach raises doubts as to whether their symmetries can be understood as weak versions of Hopf algebras–a notion and a connection they did not address.
The algebraic results of Tsuchiya and Kanie were developed into the operator algebraic approach to conformal field theory framework by A. Wassermann [98]. Utilizing von Neumann algebras, Wassermann constructed the algebras of observables for chiral model of the CFT describing the WZW model, known as the loop group conformal net.
Subsequently, the theory developed toward the direct computation and classification of the tensor categories of localized endomorphisms. By largely leaving the quantum gauge group viewpoint aside, this structural shift proved successful for 2D classifications. However, it also resulted in a lack of a unifying mathematical language between low-dimensional braided theories and higher-dimensional symmetric theories governed by standard compact gauge groups.
1.5. Our Approach to the two Problems: Quantum Gauge Groups and Drinfeld Twists
Our approach to Huang’s problem and to Doplicher-Roberts problems is derived from synthesizing the developments in affine Lie algebras by Wassermann and Toledano-Laredo on the loop group conformal net [94] with the unitary properties of quantum group fusion categories previously highlighted by Wenzl [99], from the view point of quantum symmetries in algebraic conformal field theory. In our approach primary fields play a prominent role to establish the equivalence, as originally envisaged by Mack and Schomerus. In our approach, a weak quasi-Hopf algebra is naturally associated to the conformal field theory, while a weak Hopf algebra, in a suitable new sense, is naturally associated to the quantum group fusion category, and we connect them by a Drinfeld twist.
Specifically, we solve Huang’s problem and advance Mack-Schomerus quantum symmetry program, by introducing a global quantum gauge group approach. We construct finite-dimensional weak Hopf $C^{*}$-algebras $A_{W}({\mathfrak{g}},q)$ with unitary coboundary structures associated to the fusion category of quantum groups at roots of unity. To achieve this, we utilize a specific linear fiber functor $W$ originating from the work of Wenzl on ${\mathcal{C}}({\mathfrak{g}},q)$ to the category of Hilbert spaces [99]. By applying an explicit Drinfeld twist–analogous to Drinfeld’s original proof of the Drinfeld-Kohno theorem–we implement a unitary weak quasi-tensor structure directly onto the Zhu algebra. This explicit construction allows us to naturally and directly transport the rigid, modular, and unitary tensor structures from the quantum group category to the representation category of the affine vertex operator algebra (VOA) at positive integer levels, thereby providing the desired direct link in Huang’s problem.
Our work has roots in our previous works in the setting of compact quantum groups of Woronowicz, [11], [13], [83][89]; however the different nature of categories involved here necessitates a very different framework. Woronowicz developed compact quantum groups as noncommutative function algebras in the setting of $C^{*}$-algebras via a concrete Tannaka-Krein duality for $C^{*}$-algebras extending the classical case. In his case, categories are naturally endowed with a fibre functor to the category of f.d. Hilbert spaces.
Our approach extends Woronowicz Tannaka-Krein duality on several aspects. The first aspect is that our categories do not have a fibre functor a priori. Similarly to Doplicher-Roberts setting, we must construct one naturally, compatible with the braided tensor structure of the category. We need to do this first for the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$, and then for the VOA fusion category ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$.
@@ 1.5. Our Approach to the two Problems: Quantum Gauge Groups and Drinfeld Twists
To describe our approach, we go back to the constructive approach of rigid braided tensor categories motivated by CFT, mentioned in Subsect. 1.1. We recall that Drinfeld original motivation for introducing quasi-Hopf algebras was to interpret them as symmetries to describe braided tensor categories from important models of rational conformal field theories motivated by the work by Moore and Seiberg. Although as noted in [23], they could not describe precisely simple models such as the Ising, Drinfeld evaluated quasi-Hopf algebras a good description of quantum symmetries for the WZW model in an approximate way, i.e. in the limit to infinity of the central charge.
Our approach to Huang problem is based on the program of understanding Finkelberg-Kazhdan-Lusztig combined equivalence as an extension to quantum groups of Doplicher-Roberts duality theory for the class of WZW models that are described by affine Lie algebras at positive integer levels.
@@
From this perspective, we regard the equivalence described by Drinfeld-Kohno theorem as a Tannakian counterpart that needs to be implemented on the categories appearing on the sides of Kazhdan-Lusztig-Finkelberg equivalence $\tilde{\mathcal{O}}_{\ell}\to{\mathcal{C}}({\mathfrak{g}},q)$. In building this abstract duality theory, we were inspired by Neshveyev and Tuset’s simplified proof of the Drinfeld-Kohno theorem [78] in the setting of compact quantum groups, viewing it as a Tannakian counterpart implemented at the level of Hopf and quasi-Hopf algebras. The construction of the weak Hopf algebras, weak quasi-Hopf algebras starting from the category and a linear fibre functors are not present in their work. Drinfeld original approach was the opposite, to start from the construction of Hopf and quasi-Hopf algebras. Their methods for constructing the twist are inspired by Kazhdan-Lusztig work, our methods are inspired by Drinfeld original twist. Differences between the two situations arise from the different unitary structures at roots of unity or at real values of $q$.
The weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ that we construct in [10] from the quantum group fusion category are generalizations of duals of the compact quantum groups of Wornowicz: They have algebraic and unitarity properties that extend properly that of the the duals of the compact quantum groups.
The main algebraic difference is that the coproduct $\Delta$ can not be unital in our case due to the fusion rules of the quantum grpoup fusion category. It is also not strictly associative either, but this lack of strict associativity is rather mild, because we construct a fibre functor satisfying two tensorial diagrams. Moreover the antipode satisfies the usual Hopf algebra axioms and commutes with the -involution.
The -involution of $A_{W}({\mathfrak{g}},q)$ is anticomultiplicative, rather than multiplicative as in the case of compact quantum groups. The hermitian structure of quantum group fusion categories at roots of unity was analysed in [99]. This structure implies in a non-trivial way that the adjoint bialgebra and the opposite bialgebra of $A_{W}({\mathfrak{g}},q)$ coincide, reflecting a similar structure of the hermitian structure of $U_{q}({\mathfrak{g}},q)$–although $A_{W}({\mathfrak{g}},q)$ is not obtained by a co-ideal. To form tensor product of -representations, we need a twist relating the adjoint bialgebra of $A_{W}({\mathfrak{g}},q)$ with the given bialgebra. The existence of the twist allows to unify and extend the theory with that of the compact quantum groups. Our twist is inspired by the twist that Drinfeld used for a different purpose: the proof of the Drinfeld-Kohno theorem. To construct this twist, Drinfeld made use of Drinfeld coblundary $\overline{R}$-matrix. In the case of symmetric tensor categories, the coboundary matrix is the identity. Otherwise it induces an intertwiner $\rho\otimes\sigma\to\sigma\otimes\tau$ in the fusion category, similarly to how an $R$-matrix induces the braiding as an intertwiner. However for the coboundary, the intertwiner does not satisfy the braid group relation, but has always of period $2$, extending the property of the permutation symmetry in the classical case. We summarize the structure of $A_{W}({\mathfrak{g}},q)$ by saying that this algebra is a unitary coboundary weak Hopf algebra.
1.6. Main theorem
Theorem 1.1.
Let $\mathfrak{g}$} be a complex simple Lie algebra and $k$ a positive integer, the level. Let $q=e^{\frac{i\pi}{d(k+h^{\vee})}}$, with $h^{\vee}$ the dual Coxeter number and $d$ the ratio of the square lengths of the long to the short root. Then the unitary coboundary weak Hopf algebra $A_{W}({\mathfrak{g}},q)$ induces a rigid braided tensor category structure on the module category of the affine vertex operator algebra $V_{{\mathfrak{g}}_{k}}$.
Furthermore we construct, via the Drinfeld twist method, a braided tensor equivalence between this category and the modular quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$ for all Lie types.
Regarding the comparison with the Huang-Lepowsky ribbon braided tensor structure:
  • (1)
    For the Lie types $A$, $B$, $C$, $D$, $G_{2}$ we obtain a complete identification.
  • (2)